OpenStudy (anonymous):

What is 2^3/2 equal to?

OpenStudy (anonymous):

Too many math questions get posted around here without parentheses, that desperately need to have them. A couple of other answerers here have already pointed out the ambiguity in what you wrote. You can make it clear using the additional details feature. OK, that said, it looks like you meant 2^(5/2) - 2^(3/2), because (2^5)/2 - (2^3)/2 would be so much easier. But please, in the future, practice safe math -- use parentheses! Using the rule: a^(b+c) = (a^b)(a^c), 2^(5/2) - 2^(3/2) = 2^(1 + 3/2) - 2^(3/2) = 2^(3/2) (2 - 1) = 2^(3/2) = 2^(1 + 1/2) = 2 * 2^(1/2) = 2√2

OpenStudy (anonymous):

4

OpenStudy (anonymous):

google it

OpenStudy (solomonzelman):

2^3 --- 2 or 2^3 / 2 ?

OpenStudy (jdoe0001):

\(\bf \cfrac{2^3}{2}?\)

OpenStudy (anonymous):

it is 4

OpenStudy (anonymous):

That's not one of the answer choices

OpenStudy (anonymous):

Let me post the answer choices, give me a second

OpenStudy (solomonzelman):

You didn't say if it is \(\normalsize\color{blue}{ 2^{3/2}}\) or \(\normalsize\color{blue}{ 2^{3}/2}\)

OpenStudy (jdoe0001):

\(\Large a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\)

OpenStudy (anonymous):

It is 2^(3/2)

OpenStudy (solomonzelman):

then use the formula jdoe0001 posted

OpenStudy (anonymous):

It's not one of the answers

OpenStudy (solomonzelman):

2^(3/2) = ²√ 2³ = √8

OpenStudy (solomonzelman):

Still not one of the choices ?

OpenStudy (anonymous):

It would be 2 to the square root of 8?

OpenStudy (anonymous):

\[2\sqrt{8} ?\]

OpenStudy (solomonzelman):

It is \(\normalsize\color{black}{ \sqrt[2]{8}}\), which is the exact same thing as \(\normalsize\color{black}{ \sqrt{8}}\)

OpenStudy (anonymous):

Okay, thanks for showing me how to do it in the process :D

OpenStudy (anonymous):

You never fail to disappoint

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

Did I say that wrong

OpenStudy (anonymous):

You never fail to solve the question!

OpenStudy (solomonzelman):

you said, \(\normalsize\color{black}{ 2\sqrt{8}}\) and it is not that, it is \(\normalsize\color{black}{ \sqrt[2]{8}}\)

OpenStudy (solomonzelman):

it is not \(\normalsize\color{black}{ 2\times\sqrt{8}}\), just \(\normalsize\color{black}{ \sqrt{8}}\)

OpenStudy (anonymous):

its d

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

I gave you the answer, didn't I ?